OpenStudy (anonymous):

If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value of x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

f(3)means value of function at x= 3

OpenStudy (gorv):

and function here is x-3

OpenStudy (anonymous):

so what does that mean?

OpenStudy (gorv):

that means f(x)=x-3 at x= 3 f(3)=3-3=0 which means value of x-3 at x =3

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (gorv):

A